A person jogs eight complete laps around a 400-mm track in a total time of 14.0 minmin .

Part A

Part complete

Calculate the average speed, in m/sm/s.

v¯v¯ =

3.81

m/sm/s

SubmitPrevious Answers

Answer Requested

Part B

Calculate the average velocity, in m/sm/s.
Answer #1

formula for average speed

v avg =total distance / time = 8(400 m)/14 ( 60 s) = 3.81 m/s

(B)

avergae velocity = total displacement / time = 0/14(60) = 0